A pizza restaurant manager claims that the average home delivery time for their pizza is no more than 25 minutes. A random sample of 49 home delivery pizzas was collected. The sample mean was found to be 26.5 minutes and the standard deviation was found to be 4 minutes. Is there evidence to reject the manager’s claim at alpha =.05? Perform an appropriate hypothesis test, showing the necessary calculations and/or explaining the process used to obtain the results.
Answer by stanbon(75887)

A pizza restaurant manager claims that the average home delivery time for their pizza is no more than 25 minutes. A random sample of 49 home delivery pizzas was collected. The sample mean was found to be 26.5 minutes and the standard deviation was found to be 4 minutes. Is there evidence to reject the manager’s claim at alpha =.05? Perform an appropriate hypothesis test, showing the necessary calculations and/or explaining the process used to obtain the results.
-----------------------------------
Ho: u <= 25
Ha: u > 25 (claim)
--------------
z(26.5) = (26.5-25)/4 = 0.3750
---
p-value = P(z > 0.3750 = 0.3538
----
Since the p-value is greater than 5%, fail to reject Ho.
Conclusion: The test results do not support the claim.
============
Cheers,
Stan H.